OpenStudy (kvng_leech):

whats the square root 137

rebeccaxhawaii (rebeccaxhawaii):

http://prntscr.com/bcio6r click that, then type in 13.

rebeccaxhawaii (rebeccaxhawaii):

137*

OpenStudy (kittyloli):

11.704

OpenStudy (dustin_whitelock):

Are you asking for the method of answering using long division?

OpenStudy (kittyloli):

\[\sqrt{137}=11.704\]

OpenStudy (zzr0ck3r):

The only way to write \(\sqrt{137}\) is \(\sqrt{137}\). \(137\) is prime so any other way you write it is an approximation.

OpenStudy (plainntall):

If you want to guess you can use Newton's method. f(x)=x^2-137 f'(x)=2x x-f(x)/f'(x)=x -(x^2-137)/2x so if we start with 11 since 137 is between 11^2=121 and 12^2=144 11-(121-137)/2(11) = 11 - (-16/22) = 11+ .7 = 11.7 running the approximation a second time with the value 11.7 we get 11.7 -( 136.89-137)/2(11.7) = 11.704 and again 11.704 - (136.9836-137)/2(11.704)=11.7046 you can keep doing that for as far as you want to take the decimal place using the equation x-(x^2-n)/2x) where x is your guess and n is the number you are trying to find the root for.